OpenStudy (anonymous):

Solve the equation by completing the square: x2 - 8x + 15 = 0 Choose one answer. a. x =-3 or x = 5 b. x =3 or x = -5 c. x = + 5 d. x =3 or x =5

OpenStudy (anonymous):

\[x^{2}-(a+b)x+(a*b)\] 8 = a+b 15 = a*b What are the factors of 15?

OpenStudy (anonymous):

3 and 5 rite?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thats my answer?